### v3

# Perspective: Interactions mediated by atoms, photons, electrons, and excitons

Abstract
Interactions between quasiparticles mediated by a surrounding environment are ubiquitous and lead to a range of important effects from collective modes of low temperature quantum gases, superconductivity, to the interaction between elementary particles at high energies. This perspective article is motivated by experimental progress in the fields of quantum degenerate atomic gases, cavity QED, and two-dimensional (2D) semi-conductors, which enable a systematic exploration of mediated interactions in new settings and regimes. We first describe how to microscopically calculate the quasiparticle interaction using perturbation theory, diagrammatics, and the path integral, highlighting the key role played by the quantum statistics of the quasiparticles. Recent theoretical and experimental insights into quasiparticle and mediated interactions in general obtained from atomic gases are then discussed, after which we focus on hybrid light-atom systems where a remarkable long range photon mediated interaction can be realised. Next, we describe new and puzzling results regarding the interaction between quasiparticles in 2D semiconductors. We then discuss how mediated interactions open up ways to realise new quantum phases in atomic and hybrid atom-photon systems as well as 2D semiconductors, and the perspective ends by posing some open questions and outlook.
I Introduction
An intrinsic property of quasiparticles is that they interact with each other, because the effects of one quasiparticle on its surrounding are felt by another as illustrated in Fig. 1. Such interactions lead to a range of remarkable and important effects including superconductivity caused by an attractive phonon mediated interaction between electrons that overwhelms their Coulomb repulsion Bardeen et al. (1957a, b); Schrieffer (1983), and giant magnetoresistance Grünberg et al. (1986); Parkin et al. (1990). Another prominent example of mediated interactions is that between 3He atoms in liquid 4He Bardeen et al. (1967). Also, it has been conjectured that interactions between electrons mediated by spin fluctuations are the mechanism behind high-temperature superconductivity Bednorz and Muller (1986); Kresin and Wolf (2009), and at the fundamental level all interactions between elementary particles are mediated by gauge bosons Weinberg (1995); ’t Hooft (1980); Leike (1999).
Figure 1: Cartoon of mediated interactions. Two impurities (yellow balls) perturb their surroundings, which therefore mediates an effective interaction between them.
This perspective is motivated by advances in different experimental platforms providing new possibilities to explore mediated interactions in quantum many-body systems. The first platform consists of atomic gases cooled down to very low temperatures such that their properties are governed by quantum mechanics. This realizes highly flexible many-body systems where fundamental properties such as the interaction strenght, spatial dimension, and quantum statistics of the particles can be experimentally controlled Bloch et al. (2008); Giorgini et al. (2008); Gross and Bloch (2017). By inserting such atomic gases in optical cavities, one can moreover realise entirely new hybrid light-matter systems with several exotic properties Mivehvar et al. (2021). Likewise, atomically thin transition metal dichacogenides (TMDs) like MoSe2, MoS2, WSe2, and WS2 open up new possibilities to explore truly 2D many-body physics in a solid state setting. They are direct band gap semiconductors with a rich set of spin-valley degrees of freedom, which has led to the realization of several exotic phases of matter Regan et al. (2022); Kennes et al. (2021), such as as Mott insulators Shimazaki et al. (2021), Wigner crystals Regan et al. (2020); Smoleński et al. (2021), topological superconductivity Kezilebieke et al. (2020), and the integer and fractional quantum Hall effects Li et al. (2021); Xu et al. (2023).
In this perspective, we will discuss how these versatile experimental platforms, which seem quite different but actually share many similarities, offer new ways to explore mediated interactions systematically and in regimes never realised before. This has led to new theoretical and experimental understandings as well as opened up new frontiers. The perspective article starts by setting up the basic framework for describing interactions between quasiparticles, and we then explain different theoretical approaches for calculating them, from perturbation theory over diagrammatic theory to the path integral. This highlights the key role played by the quantum statistics of the quasiparticles in determining the sign of the interaction, and also the relation between quasiparticle and general mediated interactions. We then discus theoretical and experimental insights regarding interactions mediated in atomic Fermi gases and Bose-Einstein condensates (BECs). The striking realisation of essentially infinite range interactions between atoms mediated by photons in an optical cavity are then described using both a diagrammatic and an equations of motion approach, after which we turn our attention to the experimental and theoretical advances regarding mediated interactions in TMDs. We end this perspective by summarising results and discussing some open questions and frontiers.
This article is organized as follows. The fundamental theoretical framework for describing quasiparticles and their interactions is described in Sec.II. We then explain three different ways to calculate quasiparticle interactions: perturbation theory rigorously valid for weak coupling in Sec.III, a general diagrammatic approach capable of going beying weak coupling in Sec.IV.1, and finally in Sec.IV.2 a path integral approach where mediated interactions emerge quite naturally. The concepts of the quasiparticle and mediated interactions are disentangled in Sec. V, and in Sec. VI we discuss the recent experimental and theoretical progress regarding fermion-mediated interactions in atomic quantum gases. The case of boson-mediated interactions in atomic gases is discussed in Sec. VII, whereas Sec. VII.3 is devoted to the long range interactions between atoms mediated by cavity photons. In Sec. VIII, we discuss mediated interactions between excitons in TMDs, and the emergence of collective phases due to mediated interactions are outlined in Sec. IX. We finally end with conclusions and outlook in Sec. X.
II Interactions between quasiparticles
In this section, we will set up the general framework for analysing the interactions between quasiparticles as first put forward by Landau, and which is a pinnacle in theoretical physics allowing a relatively simple yet powerful description of otherwise complicated interaction quantum many-body systems. The premise of Landau’s theory is that the energy of a many-body system can be expanded as a power series in the occupation numbers of its fundamental excitations Baym and Pethick (2008), which behave as either bosonic or fermionic particle-like entities, so-called quasiparticles, with renormalised properties, finite life-times. Keeping terms up to second order in the occupation numbers, the energy of the system is written as
$E=E_{0}+\sum_{\mathbf{p},\sigma}\varepsilon^{0}_{\mathbf{p}\sigma}n_{\mathbf{p}\sigma}+\frac{1}{2}\sum_{\mathbf{p},\mathbf{p}^{\prime},\sigma,\sigma^{\prime}}f_{\mathbf{p}\sigma,\mathbf{p}^{\prime}\sigma^{\prime}}n_{\mathbf{p}\sigma}n_{\mathbf{p}^{\prime}\sigma^{\prime}}~{},$ (1)
where $E_{0}$ is the energy of the system in the absence of any quasiparticle excitations, $\varepsilon^{0}_{\mathbf{p}\sigma}$ is the energy of a single quasiparticle with momentum $\mathbf{p}$, $n_{\mathbf{p}\sigma}$ is its occupation, and $\sigma$ denotes a possible internal quantum state such as spin. We have here and in the rest of the paper assumed a unit system volume. The first derivative of Eq. (1) gives the energy of a quasiparticle
$\varepsilon_{\mathbf{p}\sigma}=\frac{\delta E}{\delta n_{\mathbf{p}\sigma}}=\varepsilon_{\mathbf{p}\sigma}^{0}+\sum_{\mathbf{p}^{\prime}\sigma^{\prime}}f_{\mathbf{p}\sigma,\mathbf{p}^{\prime}\sigma^{\prime}}n_{\mathbf{p}^{\prime}\sigma^{\prime}}~{},$ (2)
where the second term is its energy shift due to the interactions with other quasiparticles. The second derivative gives the interaction between two quasiparticles in quantum states $\mathbf{p}\sigma$ and $\mathbf{p}^{\prime}\sigma^{\prime}$,
$f_{\mathbf{p}\sigma,\mathbf{p}^{\prime}\sigma^{\prime}}=\frac{\delta^{2}E}{\delta n_{\mathbf{p}\sigma}\delta n_{\mathbf{p}^{\prime}\sigma^{\prime}}}.$ (3)
Importantly, this interaction is an inherent property of quasiparticles in a medium since one quasiparticle affects its surroundings, which is felt by another quasiparticle. Put another way, two quasiparticles interact by exchanging modulations in their surrounding medium. For a thorough discussion of Landau’s quasiparticle theory, we refer to Refs. Baym and Pethick (2008); Fetter and Walecka (1971); Lifshitz and Pitaevskii (1980).
III Perturbation theory
We present in this section two cases where the quasiparticle interaction given by Eq. (3) can be calculated rigorously using perturbation theory.
III.1 Fermion mediated interaction
Consider $N_{a}$ mobile ”impurity” particles of mass $m_{a}$ immersed in an ideal spin-polarized (single component) gas of fermionic particles with mass $m_{b}$ and density $n_{F}$. The Hamiltonian is $\hat{H}=\hat{H}_{0}+\hat{H}_{\text{int}}$ with
$\hat{H}_{0}=$ $\sum_{\mathbf{p}}\epsilon_{\mathbf{p}a}\hat{a}^{\dagger}_{\mathbf{p}}\hat{a}_{\mathbf{p}}+\sum_{\mathbf{p}}\epsilon_{\mathbf{p}b}\hat{b}^{\dagger}_{\mathbf{p}}\hat{b}_{\mathbf{p}}~{},$ (4)
$\hat{H}_{\text{int}}=$ $g\sum_{\mathbf{p},\mathbf{p}^{\prime},\mathbf{q}}\hat{a}^{\dagger}_{\mathbf{p}+\mathbf{q}}\hat{b}^{\dagger}_{\mathbf{p}^{\prime}-\mathbf{q}}\hat{b}_{\mathbf{p}^{\prime}}\hat{a}_{\mathbf{p}}.$ (5)
Here, $\hat{a}^{\dagger}_{\mathbf{p}}$ creates an impurity particle with momentum $\mathbf{p}$ and kinetic energy $\epsilon_{\mathbf{p}a}=p^{2}/2m_{a}$, $\hat{b}^{\dagger}_{\mathbf{p}}$ creates a fermion with momentum $\mathbf{p}$ and kinetic energy $\epsilon_{\mathbf{p}b}=p^{2}/2m_{b}$, and $g$ is the interaction between the fermions and the impurity particles, which we take to be momentum independent (infinitely short range), whereas the fermions do not interact amongst themselves. While such a system may seem oversimplified at first, it has in fact been created experimentally by several groups using cold atomic gases and it features several interesting physical effects Schirotzek et al. (2009); Nascimbène et al. (2009); Kohstall et al. (2012); Koschorreck et al. (2012); Massignan et al. (2014); Cetina et al. (2016); Scazza et al. (2017); Schmidt et al. (2018); Fritsche et al. (2021). In particular, the interaction between the impurity particles and the Fermi gas leads to the formation of quasiparticles denoted as Fermi polarons, which consist of a impurity particle surrounded by a cloud of fermions.
Before we delve into the calculations, we note that while Eq. 5 relates to the interaction between a bare impurity and a fermion from the Fermi sea, Landau’s quasiparticle relates to an effective interaction between quasiparticles, which we now calculate. We begin with the interaction between two Fermi polarons assuming the interaction $g$ is weak so that perturbation theory applies. An eigenstate for no interactions ($g=0$) is $|\psi_{0}\rangle=\prod_{\mathbf{p}}\hat{a}^{\dagger}_{\mathbf{p}}\prod_{|\mathbf{k}|<k_{F}}\hat{b}^{\dagger}_{\mathbf{k}}|0\rangle$ where $|0\rangle$ is the vacuum state and $k_{F}$ is the Fermi momentum. The energy shift to first order is $E_{1}=\langle\psi_{0}|\hat{H}_{\text{int}}|\psi_{0}\rangle=N_{a}gn_{F}$, which is simply the mean-field energy coming from the density $n_{F}$ of the fermions. The second order energy shift is $E_{2}=\sum_{n\neq 0}\left|\langle\psi_{0}|\hat{H}_{\text{int}}|\psi_{n}\rangle\right|^{2}/(E_{0}-E_{n})$. The relevant unperturbed eigenstates are $|\psi_{n}\rangle=[n_{\mathbf{p}a}(n_{\mathbf{p}-\mathbf{q}a}+1)]^{-1/2}\hat{a}^{\dagger}_{\mathbf{p}-\mathbf{q}}\hat{a}_{\mathbf{p}}\hat{b}^{\dagger}_{\mathbf{k}+\mathbf{q}}\hat{b}_{\mathbf{k}}|\psi_{0}\rangle$ where $n_{\mathbf{p}a}$ is the occupation function of the impurities. This straightforwardly gives
$E_{2}=g^{2}\sum_{\mathbf{p},\mathbf{k},\mathbf{q}}\frac{n_{\mathbf{k}b}(1-n_{\mathbf{k}+\mathbf{q}b})n_{\mathbf{p}a}(1\pm n_{\mathbf{p}-\mathbf{q}a})}{\epsilon_{\mathbf{k}b}+\epsilon_{\mathbf{p}a}-\epsilon_{\mathbf{k}+\mathbf{q}b}-\epsilon_{\mathbf{p}-\mathbf{q}a}},$ (6)
where $n_{\mathbf{k}b}=1/\{\exp[(\epsilon_{\mathbf{k}b}-\mu_{b})/T]+1\}^{-1}$ is the occupation function for the fermions with chemical potential $\mu_{b}$. The upper/lower sign in the second bracket in Eq. (6) is for bosonic/fermionic impurities and originates from Bose stimulation/Fermi blocking of the scattering processes. Mathematically, it arises from the (anti)-commutation rules for (fermions) and bosons.
From Eqs. (3) and (6), we then get Yu and Pethick (2012)
$f_{\mathbf{p},\mathbf{p}^{\prime}}=\pm g^{2}\chi(\mathbf{p}-\mathbf{p}^{\prime},\epsilon_{\mathbf{p}a}-\epsilon_{\mathbf{p}^{\prime}a}),$ (7)
for the quasiparticle interaction between two Fermi polarons with momenta $\mathbf{p}$ and $\mathbf{p}^{\prime}$ to second order in $g$. Here
$\chi(\mathbf{p},\omega)=\sum_{\mathbf{k}}\frac{n_{\mathbf{k}b}-n_{\mathbf{k}+\mathbf{p}b}}{\omega+\epsilon_{\mathbf{k}b}-\epsilon_{\mathbf{k}+\mathbf{p}b}},$ (8)
is the density-density response function of an ideal Fermi gas (the Lindhard function) describing how compressible it is Giuliani and Vignale (2005) and we remind the reader that the upper/lower sign is for bosonic/fermionic impurities. In the limit of small momentum exchange and zero temperature, the quasiparticle interaction becomes
$\lim_{|\mathbf{p}|\rightarrow|\mathbf{p}^{\prime}|}f_{\mathbf{p},\mathbf{p}^{\prime}}=\mp g^{2}\mathcal{N}(\epsilon_{F})=\mp\frac{(\Delta N)^{2}}{\mathcal{N}(\epsilon_{F})},$ (9)
where $\mathcal{N}(\epsilon_{F})$ is the density of single-particle states at the Fermi energy $\epsilon_{F}$ Yu and Pethick (2012); Yu et al. (2010); Mora and Chevy (2010); Giraud and Combescot (2012). These expressions are general for any dimensionality of the system. In particular, for 3D, $\mathcal{N}(\epsilon_{F})=m_{c}k_{F}/2\pi^{2}$. In the second equality of Eq. 9, we have expressed the interaction as $g=(\partial\mu_{2}/\partial n_{1})_{n_{2}}$, which can be related to the number $\Delta N$ of fermions attracted/repelled by the impurity Yu and Pethick (2012); Yu et al. (2010). Equation (9) holds as long as one can perform perturbation theory in the interaction between Fermi polaron and its surroundings.
This derivation explicitly demonstrates the crucial role played by the quantum statistics of the quasiparticles: for fermionic impurities the quasiparticle interaction is repulsive whereas it is attractive for bosonic impurities. The origin of this sign difference is the Fermi blocking or Bose stimulation of impurity scattering on the fermions due to the presence of other impurities.
This calculation also illustrates how it the interaction between two quasiparticles is mediated by the exchange of particle-hole excitations in the Fermi sea.
III.2 Boson mediated interaction
As a second example where rigorous results are possible, we consider $N_{a}$ mobile impurities in a weakly interacting Bose gas forming a Bose-Einstein condensate (BEC) at zero temperature. Using Bogoliubov theory to describe the BEC gives the Hamiltonian
$\hat{H}=\sum_{\mathbf{p}}(\epsilon_{\mathbf{p}a}+n_{0}g)\hat{a}^{\dagger}_{\mathbf{p}}\hat{a}_{\mathbf{p}}+\sum_{\mathbf{p}}E_{\mathbf{p}}\hat{\gamma}^{\dagger}_{\mathbf{p}}\hat{\gamma}_{\mathbf{p}}+$
$+g\sum_{\mathbf{p},\mathbf{q}}\hat{a}^{\dagger}_{\mathbf{p}+\mathbf{q}}\hat{a}_{\mathbf{p}}\left[\sqrt{\frac{n_{0}\epsilon_{\mathbf{q}b}}{E_{\mathbf{q}}}}(\hat{\gamma}^{\dagger}_{-\mathbf{q}}+\hat{\gamma}_{\mathbf{q}})+\sum_{\mathbf{k}}\hat{b}^{\dagger}_{\mathbf{k}-\mathbf{q}}\hat{b}_{\mathbf{k}}\right],$ (10)
where $\hat{\gamma}_{\mathbf{p}}^{\dagger}=u_{\mathbf{p}}{\hat{b}}_{\mathbf{b}}^{\dagger}+v_{\mathbf{p}}{\hat{b}}_{-\mathbf{p}}$ creates a Bogoliubov excitation in the BEC with energy $E_{\mathbf{p}}=\sqrt{\epsilon_{\mathbf{p}b}(\epsilon_{\mathbf{p}b}+2n_{0}g_{b})}$, and $u^{2}_{\mathbf{p}}=1+v^{2}_{\mathbf{p}}=[(\epsilon_{\mathbf{p}b}+n_{0}g_{b})/E_{\mathbf{p}}+1]/2$ are the usual coherence factors, and $n_{0}$ the condensate fraction. Here ${\hat{b}}_{\mathbf{p}}^{\dagger}$ creates a boson with momentum $\mathbf{p}$ and $g_{b}$ is the strength of the short range repulsion between the bosons. We have in Eq. (10) omitted any constant terms giving the energy of the BEC ground state and all momenta for the bosons are different from zero. The Bogoliubov energy is linear at small momentum, $E_{\mathbf{p}}\approx c_{s}p$ with $c_{s}=\sqrt{n_{0}g_{b}/m_{b}}$ the BEC speed of sound, which resembles the acoustic phonon dispersion in solids. The last term describes interactions between the impurities and the BEC leading to the formation of quasiparticles coined Bose polarons.
In the perturbative regime where both $g$ and $g_{b}$ are small, one can neglect the last quartic term in Eq. (10) since its contribution to the quasiparticle interaction is suppressed by the small depletion of the BEC. For instance, in 3D, this suppression is proportional to $(n_{0}a_{b}^{1/3})^{1/2}$ Camacho-Guardian and Bruun (2018), where $a_{b}$ is the boson-boson scattering length with $g_{b}=4\pi a_{b}/m_{b}$. The interaction can now be calculated using second order perturbation theory following the same steps as in Sec III.1.
This gives the second order energy shift Yu and Pethick (2012)
$E_{2}=g^{2}\sum_{\mathbf{p},\mathbf{q}}\frac{n_{0}\epsilon_{\mathbf{q}b}}{E_{\mathbf{q}}}$ $\left[\frac{(1+n_{\mathbf{q}b})n_{\mathbf{p}a}(1\pm n_{\mathbf{p}-\mathbf{q}a})}{\epsilon_{\mathbf{p}a}-\epsilon_{\mathbf{p}-\mathbf{q}a}-E_{\mathbf{q}}}\right.$
$\left.+\frac{n_{-\mathbf{q}b}n_{\mathbf{p}a}(1\pm n_{\mathbf{p}-\mathbf{q}a})}{\epsilon_{\mathbf{p}a}+E_{-\mathbf{q}}-\epsilon_{\mathbf{p}-\mathbf{q}a}}\right],$ (11)
where $n_{\mathbf{q}b}=1/[\exp(E_{\mathbf{q}}/T)-1]^{-1}$ is the occupation function for the bosons and the upper/lower sign again is for bosonic/fermionic impurities.
Equations (3) and III.2 then again gives Eq. (7) for the quasiparticle interaction, but now with
$\chi(\mathbf{p},\omega)=\frac{2n_{0}\epsilon_{\mathbf{p}b}}{\omega^{2}-E_{\mathbf{p}}^{2}}$ (12)
the density-density response function of a weakly interacting BEC Yu and Pethick (2012).
In the limit of small momentum transfer, we obtain
$\lim_{|\mathbf{p}|\rightarrow|\mathbf{p}^{\prime}|}f_{\mathbf{p},\mathbf{p}^{\prime}}=\mp\frac{g^{2}}{g_{b}}.$ (13)
Physically, the $1/g_{b}$ dependence in Eq. (13) reflects that the interaction increases with the compressibility of the BEC. The second-order contribution in Eq. III.2 arising from the quartic term in Eq. 10 is suppressed by the gas factor $n_{0}a^{3}\ll 1$  Levinsen et al. (2017) and gives a small correction to the induced interaction, except for very low momenta where it gives an unphysical divergence for $\lim_{|\mathbf{p}|\rightarrow|\mathbf{p}^{\prime}|}f_{\mathbf{p},\mathbf{p}^{\prime}}.$
Like for the fermionic case discussed in Sec. III.1, this derivation explicitly demonstrates the key role of quantum statistics making the quasiparticle interaction attractive/repulsive for bosonic/fermionic quasiparticles.
IV Field theory
Having derived rigorous expressions for the interaction between quasiparticles mediated by a Fermi or Bose gas in the perturbative regime and demonstrated the key role of quantum statistics, we now discuss how this quasiparticle interaction is calculated in a field theoretical approach. This provides a framework to at least formally describe the regime of strong interactions and to introduce the induced interaction in general settings beyond Landau’s quasiparticle interaction.
IV.1 Green’s function formalism
Green’s functions provide a compact way to analyse interacting many-particle systems Fetter and Walecka (1971); Abrikosov et al. (2012). Consider again the $N_{a}$ impurity particles of mass $m_{a}$ in a medium of majority particles. The retarded Green’s function of an impurity with momentum $\mathbf{p}$ is given by $G_{a}(\mathbf{p},\omega)$ and it is governed by the Dyson’s equation $G_{a}^{-1}(\mathbf{p},\omega)=G^{-1}_{a0}(\mathbf{p},\omega)-\Sigma_{a}(\mathbf{p},\omega)$ where $G^{-1}_{a0}(\mathbf{p},\omega)=\omega-\epsilon_{\mathbf{p}a}$ is the noninteracting Green’s function and $\Sigma_{a}(\mathbf{p},\omega)$ the self-energy of the impurity atoms. The energy of the quasiparticle state of an impurity with momentum $\bf p$ is determined by the self-consistent equation Fetter and Walecka (1971); Abrikosov et al. (2012)
$\varepsilon_{\mathbf{p}}=\epsilon_{\mathbf{p}a}+\text{Re}\Sigma_{a}(\mathbf{p},\varepsilon_{\mathbf{p}}).$ (14)
Comparing with Eq. (2), we see in order to extract the quasiparticle interaction, we must calculate how the self-energy depends on the impurity concentration. Explicitly, we obtain from Eqs. (2) and (14)
$f(\mathbf{p},\mathbf{p}^{\prime})=\frac{\delta\varepsilon_{\bf p}}{\delta n_{\bf p^{\prime}}}=Z_{\mathbf{p}}\frac{\delta\text{Re}\Sigma_{a}(\mathbf{p},\varepsilon_{\bf p})}{\delta n_{\mathbf{p}^{\prime}}},$ (15)
where
$Z_{\mathbf{p}}=\left.\left[1-\frac{\partial\text{Re}\Sigma_{a}(\mathbf{p},\omega)}{\partial\omega}\right]^{-1}\right|_{\omega=\varepsilon_{\mathbf{p}}},$ (16)
is the quasiparticle residue and we have ignored the spin index $\sigma$ for simplicity. The presence of the residue reflects that it is only the quasiparticle part of the Green’s function that contributes to the quasiparticle interaction. Note that Eq. (15) is generally valid and not restricted to the perturbative regime. We now give two examples for calculating the self-energy and the resulting effective interaction from Eq. (15).
First, consider again the case described in Sec. III.1, i.e. mobile impurity particles in a single component Fermi gas described by the Hamiltonian in Eqs. (4)-(5). The self-energy of the impurities is simply given by the mean-field shift $\Sigma_{1a}=gn_{F}$ to first order in $g$, which is independent of the impurity concentration. The second order self-energy shown in Fig. 2(a) is
$\Sigma_{2a}({\bf p},i\omega_{p})=\pm g^{2}\int\!\frac{d^{3}p^{\prime}}{(2\pi)^{3}}\chi({\bf p}-{\bf p}^{\prime},\omega)n_{{\bf p}^{\prime}a},$ (17)
where the upper/lower sign is for bosonic/fermionic impurities with the distribution function $n_{{\bf p}a}$ and we have ignored terms independent on the impurity concentration. Taking the derivative of $\Sigma_{2a}$ according to Eq. (15) then recovers Eqs. (7)-(8) for the Landau quasiparticle interaction between two quasiparticles mediated by a Fermi gas. As illustrated in Fig. 2, taking the derivative corresponds diagrammatically to cutting the impurity line inside the self-energy. Here, the blue/red lines correspond to impurity/fermion propagators, whereas the wavy lines are the direct impurity-boson interaction. This calculation also shows that the sign difference between the effective interaction for fermonic and bosonic impurities can be understood from the fact that the mediated interaction originates from an exchange (Fock) diagram Yu and Pethick (2012); Mora and Chevy (2010); Yu et al. (2010).
The power of the Green’s function method is that provides a formalism to analyse the strong coupling regime beyond perturbation theory. An example of this is shown in Fig. 2(b), where the so-called ladder diagrams are included to infinite order in the self-energy. This approach, which describes two-body correlations between the impurities and the fermions exactly, is known to be remarkably accurate for mobile impurities interacting with a Fermi gas via a short range potential, i.e. precisely the problem at hand Massignan et al. (2014). Diagrammatically, the corresponding quasiparticle interaction is obtained by cutting one of the internal impurity lines in the self-energy due to the derivative in Eq. (15), which is illustrated in Fig. 2(b). As we shall discuss in Sec. VI(b), the resulting quasiparticle interaction agrees well with experimental results for the interaction between Fermi polarons in an atomic gas.
As a second example, consider again impurities in a weakly interacting BEC as described in Sec. III.2. The first order diagram gives the mean field shift as before and the second order self-energy is shown diagrammatically in Fig. 2(c). The impurity/Bogoliobov mode propagators are depicted by the blue/double red lines respectively. Condensate bosons are illustrated by the red dashed lines. Cutting the impurity line (taking the derivative in Eq. (15)), gives Eq. (17) with the compressibility of a BEC given by Eq. (12). This in turn recovers Eq. (7) for the effective interaction. Again, this perturbative result can be generalised to strong interactions using the ladder approximation as described in Sec. VII and Ref. Camacho-Guardian and Bruun (2018).
Figure 2: (a) Second-order self-energy of an impurity in a Fermi gas. The derivative in Eq. (15) corresponds to cutting the intermediate line and gives rise to the quasiparticle interaction shown on the right. The solid red/blue lines correspond to the bath/impurity Green’s function, the wavy black line depicts the impurity-fermion interaction $g$. (b) A typical higher-order process in the ladder approximation for an impurity in a Fermi gas. Cutting an impurity propagator yields the diagram for the quasiparticle interaction shown on the right. (c) Second-order self-energy for an impurity immersed in a BEC. Cutting the intermediate impurity line leads to a phonon-mediated interaction between the impurities (right). Here the red solid double lines correspond to a Bogoliubov mode propagator. The dashed lines correspond to condensate bosons.
IV.2 Path integral approach
Mediated interactions emerge naturally within the path integral formulation of quantum field theory Bijlsma et al. (2000); Stoof et al. (2009). We illustrate this by considering again the case of the interaction between impurity particles mediated by a weakly interacting BEC as described by the Hamiltonian in Eq. (10); a similar calculation for the case of a fermion mediated interaction can be found in e.g. Refs. Zheng et al. (2021, 2024). The starting point is the partition function $Z=\text{Tr}e^{-\beta({\hat{H}}-\mu{\hat{N}})}$ written as a path integral Orland and Negele (1998)
$Z=\int\!{\mathcal{D}}(\gamma^{*},\gamma,a^{*},a)e^{-S(\gamma^{*},\gamma,a^{*},a)},$ (18)
where the action is
$S=\sum_{p}\left[a^{*}_{p}(-i\omega_{p}+\epsilon_{\mathbf{p}a}+n_{0}g)a_{p}+\gamma^{*}_{p}(-i\omega_{p}+E_{\bf p})\gamma_{p}\right]$
$+g\sum_{p,q}a^{*}_{p+q}a_{p}\sqrt{\frac{n_{0}\epsilon_{\mathbf{q}b}}{E_{\mathbf{q}}}}(\gamma^{*}_{-q}+\gamma_{q}).$ (19)
Here, $\gamma_{p}$ are complex numbers in momentum/frequency space $p=({\bf p},i\omega_{p})$ with $\omega_{p}$ a bosonic Matsubara frequency describing the Bogoliubov modes in the BEC, and $a_{p}$ are complex/Grassman numbers depending on whether the impurities are bosons/fermions. We have omitted the quartic term in Eq. (10) as its contribution to the induced interaction is suppressed within Bogoliubov theory, for 3D by a factor $(n_{0}a_{b}^{1/3})^{1/2}$ as discussed in Sec. III.2 Camacho-Guardian and Bruun (2018).
Since there are only quadratic terms for the complex Bose field $\gamma_{p}$, it is straightforwardly integrated out using the Gaussian identity Orland and Negele (1998)
$\int\!{\mathcal{D}}(\gamma_{p}^{*},\gamma_{p})e^{\gamma^{*}_{p}G_{B}^{-1}(p)\gamma_{p}-J_{p}^{*}\gamma_{p}-\gamma_{p}^{*}J_{p}}\propto e^{-J_{p}^{*}G_{B}(p)J_{p}},$ (20)
where $G_{B}(p)=1/(i\omega_{p}-E_{\bf p})$ is the Green’s function for the Bogoliubov modes and $J_{p}=g\sqrt{n_{0}\epsilon_{\mathbf{p}b}/E_{\mathbf{p}}}\sum_{q}a^{*}_{p+q}a_{q}$. Using this, we get that the effective action for the impurities is
$S_{\text{eff}}=\sum_{p}a^{*}_{p}(-i\omega_{p}+\epsilon_{\mathbf{p}a}+n_{0}g)a_{p}+$
$+\frac{g^{2}}{2}\sum_{p,q,q^{\prime}}\frac{2n_{0}\epsilon_{{\bf p}b}}{(i\omega_{p})^{2}-E_{\bf p}^{2}}a^{*}_{q+p}a_{q}a^{*}_{q^{\prime}-p}a_{q^{\prime}}.$ (21)
From this we see that, after analytic continuation $i\omega_{p}\rightarrow\omega+i0^{+}$, the BEC mediates a frequency dependent interaction $g^{2}\chi(\bf p,\omega)$ (see Eq. (12)) between the impurities with the compressibility of the BEC given by Eq. (12). Calculating the second order Fock diagram with this interaction then gives Eq. (17) from which Eq. (7) for the quasiparticle interaction again is recovered. Here again, we ignore the quartic term in the impurity-boson coupling, which is expected to give small corrections.
V Quasiparticle interaction versus induced interaction
The analysis in Sec. IV shows that the induced interaction emerges naturally in many-body systems and is not restricted to the interaction between quasiparticles. This is explicit in Eq. (21), where an induced interaction $V_{\text{ind}}({\bf p},\omega)=g^{2}\chi({\bf p},\omega)$ between the impurities mediated by the BEC appears already in the effective Lagrangian. Indeed, the induced interaction is quite generally given by $V_{\text{ind}}({\bf p},\omega)=g^{2}\chi({\bf p},\omega)$ in the perturbative regime, where $\chi({\bf p},\omega)$ is the compressibility of the medium carrying the interaction. This expression can be generalised to strong interactions by considering the appropriate diagrams as exemplified in Sec. IV.1. The frequency dependence of the induced interaction reflects that excitations in the medium carrying the interaction propagate with a finite speed giving rise to retardation effects. Such retardation effects are very important for instance in the so-called strong coupling superconductors Mahan (2000).
The induced interaction and Landau’s quasiparticle interaction are connected by
$f(\mathbf{p},\mathbf{p}^{\prime})=\pm V_{\text{ind}}(\mathbf{p}-\mathbf{p}^{\prime},\varepsilon_{\mathbf{p}^{\prime}}-\varepsilon_{\mathbf{p}})$ (22)
where the upper/lower sign is for bosons/fermions.
It is illuminating to consider the induced interaction in the static limit ($\omega\rightarrow 0)$. Fourier transforming Eq. (12) to real space yields in 3D
$V_{\text{med}}(\mathbf{r})=-\frac{g^{2}n_{0}m_{B}}{\pi}\frac{e^{-\sqrt{2}r/\xi}}{r}.$ (23)
Equation (23) is the well-known Yukawa interaction between two static impurities separated by a distance $r$ in a BEC. Note that this interaction is always attractive as opposed to the quasiparticle interaction, where the sign depends on the quantum statistics of the impurities. The reason is that two static impurities separated by a distance are distinguishable so that quantum statistics is irrelevant. Physically, the attraction arises because one impurity increases/decreases the density of the BEC in its neighborhood for attractive/repulsive impurity-boson interaction, which attracts another impurity. The range of the Yukawa interaction in Eq. (23) is given by the BEC healing length $\xi=1/\sqrt{2m_{b}n_{b}g_{b}}$, which increases with decreasing $g_{b}$ reflecting that a softer (more compressible) BEC is more efficient at mediating an interaction as discussed in Sec. III.2.
Likewise, the interaction between two static impurities in a three-dimensional single component Fermi gas at zero temperature is given by
$V_{\text{med}}(\mathbf{r})=g^{2}\frac{m_{b}}{16\pi^{3}}\frac{2k_{F}r\cos(2k_{F}r)-\sin(2k_{F}r)}{r^{4}},$ (24)
within second order perturbation theory, which is the well-known Ruderman–Kittel–Kasuya–Yosida (RKKY) interaction Ruderman and Kittel (1954); Kasuya (1956); Yosida (1957).
VI Quantum Gases: fermion-mediated interactions
Having set up the general framework to describe and calculate mediated interactions including that between quasiparticles using perturbation theory, as well as many-body theory based on Feynman diagrams and the path integral, we now turn to describing the experimental and theoretical advances in recent years regarding this using quantum degenerate atomic gases as quantum simulators. A very powerful feature of atomic gases is that the impurity-medium interaction can be tuned using Feshbach resonances, which allows one to explore mediated interactions systematically and in regimes never realised before. We first discuss interactions mediated by fermions after which we turn our attention to interactions mediated by a BEC.
VI.1 Mediated interactions in 3D
Two experiments based on atomic Bose-Fermi mixtures measured a fermion mediated interaction between bosons in a BEC in the weak coupling regime. In the first experiment, the presence of a surrounding quantum degenerate 6Li gas was observed to decrease the radius of a 133Cs BEC and also give rise to a collapse instability DeSalvo et al. (2019). Due to the small mass of 6Li atoms compared to that of 133Cs atoms, the frequency dependence of the mediated interaction was expected to be weak and the results were consistent with an attractive interaction between the bosons given by the perturbative result Eq. (9). A second experiment used Ramsey spectroscopy to measure the frequency shift of a hyperfine transition for bosonic 87Rb atoms immersed in a degenerate 40K Fermi mass Edri et al. (2020). The experimental results were found to agree with a mean-field shift from a mediated interaction given by Eq. (9).
A recent experiment systematically probed the quasiparticle interaction between Fermi polarons formed by 40K or 41K atoms in a degenerate Fermi gas of 6Li atoms Baroni et al. (2024a). Using RF spectroscopy, the polaron energy was measured as a function of the quasiparticle particle concentration $n$ and compared to a momentum average of Eq. (2), i.e. $\varepsilon=\varepsilon^{0}+\bar{f}n$, since experiment had no momentum resolution. The quasiparticle interaction was explored well beyond the perturbative regime using a Feshbach resonance to tune the K-Li interaction. Moreover, by selecting either fermionic 40K or bosonic 41K atoms leaving everything else essentially unchanged, the effects of quantum statistics could be investigated directly.
Figure 3: Figures from Ref. Baroni et al. (2024a). Upper panel: The polaron energy as a function of its concentration for 41K atoms with Li-K interaction strength $X=-1/k_{F}a=0.98$. Lower panel: The polaron-polaron interaction for fermionic 40K impurity atoms (red squares) and bosonic 41K impurity atoms (green circles). The dashed lines are the perturbative result Eq. (9) and the solid lines are a strong coupling theory for the quasiparticle interaction shown diagrammatically in Fig. 2(b).
An example of such a measurement of the polaron energy as a function of its concentration is shown in the upper panel of Fig. 3 for the case of 41K impurities. The Li-K interaction is moderately strong with a scattering length $a$ giving $0.981/k_{F}a=-$ 0.98, where $k_{F}$ is the Fermi momentum of the 6Li gas. As can be seen, the energy of the polaron decreases with increasing concentration corresponding to an attractive average interaction between the polarons, which can be extracted from the slope. The lower panel of Fig. 3 shows the interaction $\bar{f}$ obtained by carefully performing such measurements for different scattering lengths and concentrations, both for 40K (red points) and 41K (green points) impurities. The first thing to notice is that the quasiparticle interaction is repulsive/attractive for fermionic/bosonic impurities thereby showing the key role of quantum statistics discussed in Sec. III. Second, the strength of the effective interaction agrees well with the second order result given by Eq. (9) (dashed lines) for weak to moderately strong Li-K interaction without any fitting parameters. The experiment thus confirms two predictions of Landau’s Fermi liquid theory: The strength of the quasiparticle interaction and its sign dependence on the quantum statistics of the quasiparticles. The perturbative result for the quasiparticle interaction however breaks down strong Li-K interactions across the Feshbach resonance where $a$ diverges as can be seen in Fig. 3. Here, a theory for the quasiparticle interaction including retardation effects and Feshbach molecule formation (solid lines) was able to explain the results for strong and attractive Li-K interactions (negative $a$). It is based on Eq. (15) together with a ladder approximation for the impurity self-energy generalised to a non-zero impurity concentration and is shown diagrammatically in Fig. 2(b). The experimental results for strong and repulsive interactions were not recovered by this theory, which may be due to the break-down of Fermi liquid theory or the formation of attractive instead of repulsive polarons, and requires further analysis.
The interaction between two static (immobile) impurities is simpler to explore than between mobile impurities, since they can be treated as scattering potentials without any recoil. This limit should correspond to an infinite impurity mass apart from the fact that the role of quantum statistics is lost since the localised impurities are distinguishable. This case is analogous to the Casimir force between two parallel conducting plates separated by a distance $r$ much smaller than the sides of the mirrors. The vacuum fluctuations of the electromagnetic field depend on the boundary conditions imposed by the plates and therefore, they are different with or without plates, leading to what is referred to as Casimir energy, scaling as $1/r^{3}$ Casimir and Polder (1948). The interaction between two static impurities mediated by an ideal Fermi gas was studied theoretically in Refs. Nishida (2009); Enss et al. (2020). By solving the scattering problem of two short range interaction potentials exactly, it was shown that interaction can be very different from the RKKY form for strong interactions due to the presence of (Efimov) states, where one fermion is bound to the two impurities, see Fig. 4. These bound states lead to resonances and sign changes in the scattering length Enss et al. (2020). For weak interactions, the perturbative RKKY interaction given by Eq. (7) was recovered.
Figure 4: Figure from Ref. Enss et al. (2020). Mediated interaction between two static impurities in a Fermi gas for different impurity fermion scattering lengths.
VI.2 Mediated interactions in 1D
Two mobile bosonic impurities in a 1D Fermi gas were analysed using the Bethe ansatz and an effective model Huber et al. (2019). A decrease in the energy was interpreted as an attractive mediated interaction between the impurities, which was shown to hold bound states in certain regimes. The properties of two fermionic impurities interacting with a bath five fermions of another kind in a 1D harmonic trap were explored using a numerical approach for solving the time-dependent many-body Schrödinger equation. This method relies on expanding the many-body wave function in terms of a time-dependent and variationally optimized basis. It permits to extract the interspecies and intraspecies correlation functions, the equilibrium features, and their dynamical properties Mistakidis et al. (2019a). The average distance between the two fermions in the trap were found to decrease with increasing impurity-bath interactions, which was interpreted as the presence of an attractive interaction mediated by the bath. Similar results were found indicative of an attractive mediated interaction for two bosonic impurities interacting with six fermions in a 1D harmonic trap Mukherjee et al. (2020).
The interaction between two static impurities separated by a distance $r$ in 1D quantum liquids was explored following a path-integral formalism Recati et al. (2005). For a non-interacting Fermi gas, the interaction between two impurities in the limit of weak fermion-impurity interaction is quadratic in the impurity-fermion interaction $\gamma=g/v_{F}$
$V(r)=-\gamma^{2}\frac{v_{F}}{2\pi r}\cos(2\pi p_{F}r),$ (25)
where $v_{F}$ is the Fermi velocity. It is mediated the exchange of a particle-hole excitations, exhibits Friedel oscillations, and is the 1D analogue of the RKKY interaction in Eq. (24). For strong impurity-fermion interaction, the mediated interaction takes the form
$V(r)=\frac{v_{F}}{2\pi r}\text{Re}\,\text{Li}_{2}(e^{2ip_{F}r}),$ (26)
where $\text{Li}_{2}$ is the di-logarithmic function. Remarkably, the mediated interaction is independent of the impurity-fermion interaction strength in this regime. When the two static impurities are immersed in an interacting Fermi gas forming a Luttinger liquid, the mediated interaction is proportional to $1/r$ for large distances Recati et al. (2005).
Interestingly, the theoretical framework used to study Fermi polarons has found direct applications in nuclear physics, such as cluster formation in neutron-rich matter Tajima et al. (2024a); Guo and Tajima (2024); Tajima et al. (2024b).
VII Quantum Gases: Boson-Mediated Interactions
In this section, we turn to the case of interactions between impurities in BECs motivated by the rapid progress in ultracold atomic gas experiments.
VII.1 Mediated interactions in 3D
As mentioned above, a very powerful feature of atomic gases is that the interaction between the impurity atoms and the surrounding majority atoms can be tuned to be very strong using a Feshbach resonance. In order to analyse this regime, a non-perturbative diagrammatic theory for the quasiparticle interaction between Bose polarons was developed in Ref. Camacho-Guardian and Bruun (2018). The theory is based on using an extended self-energy for the impurities in Eq. (15), since it turns out to be necessary to go beyond the ladder approximation to recover the second order result given by Eqs. (7) and (12) for weak interactions. The resulting quasiparticle interaction is
$f(\mathbf{p}_{1},\mathbf{p}_{2})=\pm n_{0}Z_{\mathbf{p}_{1}}Z_{\mathbf{p}_{2}}\left[\mathcal{T}^{2}(p_{1})G_{11}(p_{1}-p_{2})+\right.$
$\left.\mathcal{T}^{2}(p_{2})G_{11}(p_{2}-p_{1})+2\mathcal{T}(p_{1})\mathcal{T}(p_{2})G_{12}(p_{2}-p_{1})\right].$ (27)
Here, $n_{0}$ is the density of the BEC, $Z_{\mathbf{p}}$ is the residue of a Bose polaron with momentum $\bf p$, $p_{i}=({\mathbf{p}}_{i},\varepsilon_{\mathbf{p}_{i}})$ is the on-shell polaron momentum/energy, and $G_{11}$ and $G_{12}$ are the normal and anomalous Green’s functions of the BEC Camacho-Guardian and Bruun (2018). Equation (27) is a strong coupling extension of the second order result, which includes two-body impurity-boson correlations exactly. It can be obtained basically by substituting the weak coupling scattering matrix ${\mathcal{T}}_{v}=2\pi a/m_{r}$ with the ladder expression $\mathcal{T}(p)={\mathcal{T}}_{v}/[1-{\mathcal{T}}_{v}\Pi(p)]$ with $\Pi(p)$ the boson-impurity pair propagator.
Using this approach, it was shown that the quasiparticle interaction between polarons can be strong close to impurity-boson resonance with $01/k_{n}a=$ 0, which leads to significant energy shifts for the polaron energy. This is shown in Fig. 5, where the polaron energy is plotted a function of its concentration for a bosonic impurity in a BEC with mass ratio $m/m_{B}=1$ (left) and a fermionic impurity with mass ratio $m/m_{B}=40/7$ (right) corresponding to 40K impurity atoms in a 7Li BEC.
Figure 5: Figures taken from Ref. Camacho-Guardian and Bruun (2018). Top: The energy shift $\Delta\epsilon=\epsilon(n)-\epsilon(0)$ of a Bose polaron, where $\epsilon(n)$ is the polaron energy for concentration $n$ and $x=n/n_{B}$ with $n_{B}$ the concentration of the BEC. The mass ratio is $m/m_{B}=1$, the temperature is $T=0.4T_{C}$ with $T_{C}$ the critical temperature of the BEC, and the boson-boson scattering length is $k_{n}a_{B}=0.2$. The energy shift is shown for various impurity-boson scattering lengths $k_{n}a$ and the perturbative result $\Delta\epsilon=-ng^{2}/g_{b}$ from Eq. (13) is shown as red squares for $101/k_{n}a=-$ 10. Bottom: The energy shift of a Bose polaron for mass ratio $m/m_{B}=40/7$, $T=0$, $k_{n}a_{B}=0.2$, and various impurity-boson scattering lengths.
We see that the energy shift is negative for a bosonic impurity (top) corresponding to an attractive polaron-polaron interaction, whereas it is positive for a fermionic impurity (bottom) corresponding to a repulsive interaction consistent with the results in Sec. III.2. Even though such energy shifts should be detectable, the quasiparticle interaction between Bose polarons remains to be observed, possibly due to the generally broad experimentally energy spectra. It has been shown that an alternative way to observe the BEC mediated interaction is via the sizable frequency shift it causes for the out-of-phase dipole mode of two Fermi gases separated by a BEC as compared to their in-phase mode Suchet et al. (2017)
Using the interaction given Eq. (27) in an effective Schrödinger equation for two polarons, it was predicted to be strong enough to support bound states, i.e. bi-polarons Camacho-Guardian et al. (2018). The results for the binding energy of the bi-polaron were furthermore shown to compare well with Monte-Carlo calculations, and to recover exact results for the bound states of the Yukawa interaction given by Eq. (23) for weak interactions. The same approach was also used to predict bound states between two Bose polarons both below and above the two-particle continuum in a 2D square lattice due to an attractive quasiparticle interaction Ding et al. (2023). Within the Fröhlich model, valid for weak interactions, bipolarons were furthermore predicted using an all-coupling treatment in Ref. Casteels et al. (2013). The problem of a gas of diluite polarons has been studied recently with perturbative a field-theory and a quantum–Monte Carlo (QMC) method in Ref. Ardila (2022). Squeezing and entanglement of distinguishable polarons have been studied in the context of two quantum Brownian particles coupled to a bosonic bath Charalambous et al. (2019).
The induced interaction between two static impurities was calculated within a Gross-Pitaevskii theory (GPT) and its nonlocal extension (NLGPT) Drescher et al. (2023). In this approach, the induced interaction between the impurities is determined by comparing the energy of two impurities $E_{d}$ fixed at a distance $d$ with the energy of two individual impurities infinitely far apart $E_{\infty}$, such that $V_{\text{ind}}=E_{d}-E_{\infty}$. For short distances compared to the interparticle spacing, it was shown that the interaction is dominated by the three-body physics of a single boson bound between the two static impurities, giving rise to an Efimov-type $V_{\text{ind}}\propto 1/r^{2}$ scaling. For larger distances of the order of the coherence length of the BEC, the interaction has the Yukawa form Eq.(23). This crossover between Efimov and Yukawa scalings of the mediated interaction was found for every impurity-boson interaction strength contrary to an earlier analysis based on a variational calculation restricted to a single Bogoliubov mode, which found a Yukawa potential for weak interactions and an Efimov interaction for resonant impurity-boson interaction Naidon (2018). The different regimes of the interaction between two static impurities mediated by a BEC of density $n$ are shown in Fig. 6 as a function of the impurity separation $d$ and the boson-impurity scattering length $a_{ib}$. In Ref. Naidon (2018), the interplay between mediated interactions and three-body Efimov states were furthermore explored.
In the same spirit, the interaction between two static impurities mediated by a general superfluid with a linear phonon dispersion was calculated using effective field theory Fujii et al. (2022). Assuming weak and short-range impurity-superfluid interactions, it was shown that the mediated interaction is dominated by the exchange of two phonons for large distances $r\gg\xi$, resulting in a $1/r^{7}$ scaling instead of the Yukawa interaction mediated by one phonon exchange dominant for shorter distances. This result holds in general for a superfluid with a low energy linear phonon dispersion with the sound velocity determining the strength of the interaction.
Considering again two static impurities in a BEC, the effects of boson-boson interactions were shown to be important for the mediated interaction and bi-polaron formation for strong impurity-boson interactions Jager and Barnett (2022). The interaction between two static impurities in a BEC was also analysed using a path integral approach Panochko and Pastukhov (2022).
Figure 6: Figure from Ref. Drescher et al. (2023). Sketch of the regimes of Efimov and Yukawa scaling of the mediated interaction between two static impurities in a BEC as a function of the inverse of the impurity-boson scattering length $a_{IB}$ and the distance between the impurities. Here, $n$ is the density of the bosonic bath.
Induced interactions are not limited to the case where the impurity-boson interaction is short-ranged. For instance, progress creating hybrid ion-atom systems Grier et al. (2009); Härter et al. (2012); Kleinbach et al. (2018); Tomza et al. (2019); Dieterle et al. (2021) has stimulated much attention to charged polarons Casteels et al. (2011); Astrakharchik et al. (2021); Christensen et al. (2021, 2022); Pessoa et al. (2024) and their interactions. The interaction between two static ions immersed in a neutral atomic BEC or Fermi gas was considered in Ref. Ding et al. (2022). For the weakly interacting case, it was shown that the interaction between two ions mediated by a BEC, exhibits a Yukawa form for intermediate distances, and a $1/r^{4}$ power-law behaviour for large distances, reminiscent of the ion-atom interactions. This Yukawa interaction for long distances arises also for the mediated interaction between Rydberg impurities Wang et al. (2015). For strong atom-ion interactions, a calculation similar to that leading to Eq. 27 shows that for large distances the mediated interaction is proportional to $\mathcal{T}(0,0;0)/r^{4}$ where $\mathcal{T}(0,0;0)$ is the zero energy/momentum atom-ion scattering matrix. For two static ions interacting weakly with a Fermi gas, the mediated interaction is given by a power-law for large density and by a RKKY form for low density. Using Monte Carlo simulations, the interaction between two static impurities in a weakly interacting BEC was studied Astrakharchik et al. (2023). The ions were found to strongly perturb the BEC in their surroundings and different regimes related to the formation of bi-polarons were identified.
VII.2 Mediated interactions in 1D
The interaction between two mobile impurities mediated by bosons in a 1D harmonic trap has been studied using a numerical solution to the time-dependent many-body Schrödinger equation. By calculating the average distance between the impurities as well as non-equilibrium dynamics, signatures of attractive and repulsive mediated interactions were identified and bound bi-polaron states predicted Mistakidis et al. (2020a, 2019b, b); the (2023).
The interaction between two static impurities separated by a distance $r$ in a 1D repulsively interacting Bose gas has been studied by numerous authors and is illustrated in Fig. 7(a). For short distances $r\lesssim\xi$ compared to the healing length of the bosons, the dressing clouds of the two impurities overlap giving rise to an interaction Klein and Fleischhauer (2005); Recati et al. (2005); Dehkharghani et al. (2018)
$V(r)=-\sqrt{\frac{n_{0}m_{B}}{g_{b}}}g^{2}e^{-2r/\xi}.$ (28)
This can be derived from a Fourier transform of the 1D version of Eq. 12 and corresponds to the exchange of a single phonon between the impurities as illustrated in Fig. 7 (b). This diagram corresponds to the one described in detail using second-order perturbation theory in Fig. 2(c), where impurities exchange a Bogoliubov mode from the condensate. For large distances $r\gg\xi$ where the two dressing clouds do not overlap, the interaction between the impurities is due to quantum fluctuations in close analogy with the Casimir effect as already discussed in Sec. VI. This gives rise to the interaction Schecter and Kamenev (2014); Pavlov et al. (2019); Reichert et al. (2019, 2019)
$V(r)=-\frac{g^{2}m_{B}\xi^{3}}{32\pi r^{3}},$ (29)
exhibiting the same $1/r^{3}$ scaling as the Casimir force. It is caused by the exchange of two phonons as illustrated in Fig. 7 (c). Further studies concerning polaron interactions and bipolarons in 1D emphasised the deformation of the superfluid by the impurities Will et al. (2021); Brauneis et al. (2021), the presence of lattice 1D bipolarons Isaule et al. (2024), and bipolarons maximally entangled appearing in the form of Bell states Anh-Tai et al. (2024).
Figure 7: Figure from Ref. Schecter and Kamenev (2014). (a) The interaction between two static impurities mediated by a 1D Bose gas. (b) Single-phonon exchange responsible for the exponential decay of the interaction at short distances. Indeed, this process corresponds to the one discussed in Fig. 2(c). (c) Two-phonon exchange giving rise to a long-range Casimir-like interaction between the impurities. The solid lines correspond to the impurity propagator while the wavy line illustrate the phonon modes.
VII.3 Photon-mediated interactions
With the strong coupling of atoms to photons in optical cavities, a new way to engineer hybrid light-matter states has emerged Carusotto and Ciuti (2013); Mivehvar et al. (2021). Quantum optical lattices refer to ultracold quantum gases confined to optical lattices embedded in high-finesse cavities, where the quantum nature of the photons and the atoms have to be treated on the same footing Maschler et al. (2008). As we will now describe, the photons mediate a long-range interaction between the atoms, which can lead to a range of interesting phenomena Park et al. (2022). In this subsection, we first discuss how derive this interaction following a diagrammatic procedure as discussed in Sec. IV.1 in this Perspective. We then connect this to a more conventional derivation using the the Heisenberg equations of motion Mivehvar et al. (2021).
Figure 8: Figure from Ref. Mivehvar et al. (2021): Cartoon of $N$ atoms loaded into an external potential $V_{\text{ext}}$ and sandwiched between two mirrors (grey plates). Cavity photons mediate long-range atomic interactions between atoms in its ground state. The interplay between the tunneling dynamics, on-site atomic interactions, and induced infinite-range interactions between the atoms leads to a rich phase diagram.
VII.3.1 General Theoretical framework
To establish a general model that describes the photon mediated interaction between atoms, we consider the scheme of Fig. 8. There, an ensemble of $N$ two-level atoms having a transition frequency $\omega_{a}$ between its internal states $\{|g\rangle,|e\rangle\}$ is inside a single-mode linear cavity with wavelength $\lambda_{c}$ and resonance frequency $\omega_{c}=2\pi c/\lambda_{c}=ck_{c}$, with $c$ the speed of light. The atoms are strongly coupled to the cavity photons with an amplitude $\mathcal{G}(\mathbf{r})=\mathcal{G}_{0}\cos{k_{c}x}$, where the oscillatory factor reflects the amplitude of the cavity mode. They are also subjected to a transverse pump laser with polarization $\mathbf{\epsilon}_{p}$ and frequency $\omega_{p}$ close to the cavity resonance frequency. The coupling between this transverse pump laser and the atoms is described by the position dependent Rabi frequency $\Omega(\mathbf{r})=\Omega_{0}\cos{k_{c}y}$. In the rotating frame of the pump laser, the system is described by $\hat{H}=\hat{H}_{0}+\hat{H}_{\text{int}}$, with $\hat{H}_{0}$ given by
$\hat{H}_{0}=\sum_{n=g,e}\int{\hat{\psi}}_{n}^{\dagger}(\mathbf{r})\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\text{ext}}(\mathbf{r})\right]{\hat{\psi}}_{n}(\mathbf{r})\text{d}\mathbf{r}$
$-\hbar\Delta_{c}{\hat{a}}^{\dagger}{\hat{a}}-\hbar\Delta_{a}\int\hat{\psi}_{e}^{\dagger}(\mathbf{r})\hat{\psi}_{e}(\mathbf{r})\text{d}\mathbf{r}$
+$+\hbar\int\{\hat{\psi}_{e}^{\dagger}(\mathbf{r})\left[\Omega(\mathbf{r})+\mathcal{G}(\mathbf{r})\hat{a}\right]\hat{\psi}_{g}(\mathbf{r})+\text{H.c.}\}\text{d}\mathbf{r}.$ (30)
Here ${\hat{\psi}}_{n}(\mathbf{r})={\hat{\psi}}_{n}(\mathbf{r},t)$ and $\hat{a}=\hat{a}(t)$ are the slowly-varying atomic and photonic annihilation field operators, $V_{\text{ext}}(\mathbf{r})$ is an external trapping potential for the atoms, $\Delta_{a}=\omega_{p}-\omega_{a}$ is the difference between the pump-laser frequency and the atomic transition frequency, and $\Delta_{c}=\omega_{p}-\omega_{c}$ the relative frequency between the pump-laser and the cavity frequencies. The atomic two-body interactions are represented by $\hat{H}_{\text{int}}$, which in the limit of low temperatures and low density are described by means of a contact interaction
$\hat{H}_{\text{int}}=\frac{1}{2}\sum_{n=g,e}g_{nn}\int{\hat{\psi}}_{n}^{\dagger}(\mathbf{r}){\hat{\psi}}_{n}^{\dagger}(\mathbf{r}){\hat{\psi}}_{n}(\mathbf{r}){\hat{\psi}}_{n}(\mathbf{r})$
$+g_{eg}\int{\hat{\psi}}_{e}^{\dagger}(\mathbf{r}){\hat{\psi}}_{g}^{\dagger}(\mathbf{r}){\hat{\psi}}_{g}(\mathbf{r}){\hat{\psi}}_{e}(\mathbf{r}).$ (31)
The coefficients $g_{ee}$, $g_{gg}$, and $g_{eg}$ can be written in terms of the $s$-wave scattering lengths $a_{nn^{\prime}}$ as $g_{nn^{\prime}}=4\pi\hbar^{2}a_{nn^{\prime}}/m$. Here it must be pointed out that for fermionic atoms, owing to Pauli exclusion principle no interactions take place among same internal states. The scattering lengths $a_{nn^{\prime}}$ can be varied externally through Feshbach resonances as a magnetic field is tuned.
Figure 9: Diagrammatic representation of the photon mediated interaction. (a) Mediated interaction to lowest order in $\Delta_{a}$ involving the exchange of a single photon. The solid orange lines represent the ground state propagator, the red lines depict the excited state propagator. As illustrated in the inset, the photon exchange process arises from a second-order process where an atom from the ground state is excited via a classical pump field (crossed dot with the black wiggly line), which then emits a photon (wavy blue line), which is absorbed in the reversed process from the inset. (c) The effective interaction between ground state atoms is denoted by $\mathcal{D}.$
Diagrammatic Formulation.- Following the Green’s function formalism discussed in Sec. IV.1, we now construct the relevant diagrams for the photon-mediated interaction among ground state atoms. Such a diagrammatic approach was recently used to analyse the Sachdev–Ye–Kitaev model in a quantum cavity Uhrich et al. (2023).
First, our main assumption is that the pump and cavity frequencies are far detuned from the excited state such that $\Delta_{a}$ is large. This is the so-called dispersive regime where the population of the excited state is small and one can write their Green’s function as $G_{e}(k)\simeq 1/\Delta_{a}$. The leading diagram in $1/\Delta_{a}$ for the photon-mediated interaction between the ground state atoms can then be drawn as in Fig. 9(a), which intuitively can be interpreted from left to right as: a ground state atom (orange line) is excited to the excited state (red line) by the classical pump field $\Omega$ (red cross with the black wiggly line). The excited state then decays to the ground state again by emitting a photon into cavity mode (wavy blue line). The cavity photon is subsequently absorbed by a second ground state atom, which is then excited. This atom finally decays back into the ground state atom via stimulated emission into the classical field. This completes the mediated interaction, as two ground state atoms effectively exchange a photon. As illustrated in Fig. 9(b), the processes highlighted by the box can be summarized into an effective interaction $\eta({\mathbf{r}})=\Omega({\mathbf{r}})\mathcal{G}({\mathbf{r}})/\Delta_{a}$ between a ground state atom and a photon. Thus, as for the other mediated interactions discussed in this perspective, the photon mediated interaction between the ground state atoms is quadratic in $\eta$. Finally, assuming that the dynamics of the cavity photon can be ignored when $\Delta_{c}$ is large so that its Green’s function is $G_{c}(k)\simeq 1/\Delta_{c}$ and taking into account the two processes in Fig. 9(c) gives
$\mathcal{D}(\mathbf{r},\mathbf{r}^{\prime})=\frac{2\Omega_{0}^{2}\mathcal{G}_{0}^{2}}{\Delta_{a}^{2}\Delta_{c}}\cos(k_{c}x^{\prime})\cos(k_{c}y^{\prime})\cos(k_{c}x)\cos(k_{c}y),$ (32)
for the photon mediated interaction between the ground state atoms. As described below, one reaches this expression for $\mathcal{D}(\mathbf{r},\mathbf{r}^{\prime})$, following a standard procedure for the Heisenberg equations of motion for the atom and photon field operators. Note that this interaction is infinite range reflecting that the cavity mode is assumed to cover the whole system. The photon-mediated interactions can be derived alternatively using a path-integral formalism as in Sec. IV.2. This approach has been employed for the study of supersolid phases of bosonic atoms inside two crossed optical cavities Lang et al. (2017); Piazza et al. (2013).
Heisenberg equations of motion.- It is illuminating to present yet another derivation of the photon mediated interaction using the Heisenberg equations of motion, which is common in the field of quantum optics. To do this, we closely follow the approach in Ref. Mivehvar et al. (2021).
Let us now start from the Heisenberg equations for the atomic and photonic field operators
$i\hbar\frac{\partial{\hat{\psi}}_{e}(\mathbf{r})}{\partial t}=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\text{ext}}(\mathbf{r})-\hbar\Delta_{a}+g_{ee}{\hat{\psi}}_{e}^{\dagger}(\mathbf{r}){\hat{\psi}}_{e}(\mathbf{r})\right.+$
$\left.+g_{eg}{\hat{\psi}}_{g}^{\dagger}(\mathbf{r}){\hat{\psi}}_{g}(\mathbf{r})\right]{\hat{\psi}}_{e}(\mathbf{r})+\hbar\left[\Omega(\mathbf{r})+\mathcal{G}(\mathbf{r})\hat{a}\right]{\hat{\psi}}_{g}(\mathbf{r}),$ (33)
and
$i\hbar\frac{\partial{\hat{\psi}}_{g}(\mathbf{r})}{\partial t}=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\text{ext}}(\mathbf{r})+g_{gg}{\hat{\psi}}_{g}^{\dagger}(\mathbf{r}){\hat{\psi}}_{g}(\mathbf{r})\right.$
$\left.+g_{eg}{{\hat{\psi}}_{e}^{\dagger}(\mathbf{r}){\hat{\psi}}_{e}(\mathbf{r})}\right]{\hat{\psi}}_{g}(\mathbf{r})+\hbar\left[\Omega(\mathbf{r})+\mathcal{G}(\mathbf{r})\hat{a}^{\dagger}\right]{\hat{\psi}}_{e}(\mathbf{r})$ (34)
$i\hbar\frac{\partial\hat{a}}{\partial t}=-\hbar\Delta_{c}\hat{a}+\hbar\int\mathcal{G}(\mathbf{r}){\hat{\psi}}_{g}^{\dagger}(\mathbf{r}){\hat{\psi}}_{e}(\mathbf{r})\text{d}\mathbf{r}.$ (35)
We now use that in the dispersive regime where the pump and cavity frequencies are far detuned from the atomic transition frequency, the fastest time scale is $1/\Delta_{a}$. This means that the excited state is minimally populated and therefore has only negligible effects on the dynamics. Using a procedure known as adiabatic elimination, the steady state of the atomic $|e\rangle$ field operator becomes $\hat{\psi}_{e,ss}(\mathbf{r})\approx\frac{1}{\Delta_{a}}[\Omega(\mathbf{r})+\mathcal{G}(\mathbf{r}){\hat{a}}]\hat{\psi}_{g}(\mathbf{r})$. After substituting this and ${\hat{\psi}}_{e}^{\dagger}(\mathbf{r}){\hat{\psi}}_{e}(\mathbf{r})\propto 1/\Delta_{a}^{2}$ into the equation of motion for ${\hat{\psi}}_{g}(\mathbf{r})$, one arrives at
$i\hbar\frac{\partial{\hat{\psi}}(\mathbf{r})}{\partial t}=\left[-\frac{\hbar^{2}\nabla^{2}}{4m}+V_{\text{ext}}(\mathbf{r})+\hbar V(\mathbf{r})+\hbar U(\mathbf{r})\hat{a}^{\dagger}\hat{a}\right.$
$\left.+\hbar\eta({\mathbf{r}})(\hat{a}^{\dagger}+\hat{a})+g_{0}\hat{n}(\mathbf{r})\right]\hat{\Psi}(\mathbf{r}),$ (36)
and
[…]
Here, we have made the identifications, ${\partial_{t}{\hat{\psi}}_{g}(\mathbf{r})=\partial_{t}{\hat{\psi}}}(\mathbf{r})$, $V({\mathbf{r}})=\Omega^{2}({\mathbf{r}})/\Delta_{a}$, $U({\mathbf{r}})=\mathcal{G}({\mathbf{r}})/\Delta_{a}$, $\hat{n}(\mathbf{r})={\hat{\psi}}^{\dagger}_{g}(\mathbf{r}){\hat{\psi}}_{g}(\mathbf{r})$, and $g_{gg}=g_{0}$. The above equations establish the coupled dynamics between matter and light.
Figure 10: Figure from Ref. Landig et al. (2016): Experimental observation of cavity-induced many-body phases. Superfluid phase (red), Mott insulator (yellow), charge density wave (blue) and supersolidity (violet).
[…]
with $\mathcal{D}(\mathbf{r},\mathbf{r^{\prime}})$ given by Eq. 32.
VII.3.2 Experiments and theoretical predictions
The long range photon mediated interaction has given rise to a number of groundbreaking experiments including the observation of the Dicke phase transition in a BEC coupled with cavity modes Kristian Baumann and Esslinger (2010); Baumann et al. (2011); Brennecke et al. (2007); Klinder et al. (2015). In pioneering experiments, four different quantum phases in a bosonic gas of 87Rb atoms in layers of 2D square optical lattices in an optical cavity were observed, which were driven by the competition between the long range photon mediated interaction and a short range atomic interaction Deng and Yi (2023); Li et al. (2017); Helson et al. (2023); Landig et al. (2016); Léonard et al. (2017a). As summarized in Fig. 10, superfluid, supersolid, Mott insulating, and charge density wave phases were found in square lattices. In particular, the supersolid state is associated with the spontaneous breaking of two symmetries: the $U(1)$ symmetry related to the phase of the condensate wave function giving rise to superfluidity, and translational symmetry that produces the lattice structure. The use of ring cavities has enabled the possibility of realizing other states and self-ordered phases. Boddeda et al. (2016); Firstenberg et al. (2016); Yang et al. (2022); Ostermann et al. (2020); Suarez et al. (2022); Hao et al. (2019); Sheng et al. (2017); Vaneecloo et al. (2022).
Other interesting results include photon-mediated atomic interactions in multimode cavity QED  Vaidya et al. (2018), Higgs and Goldstone modes in a supersolid quantum gas Léonard et al. (2017b), spinor BEC self-ordering in a cavity Kroeze et al. (2018), and the formation of a spin textures in a quantum gas coupled to a cavity Landini et al. (2018). Also, disorder driven density and spin self-ordering of a Bose-Einstein condensate as well as the realisation of the Sachdev-Ye-Kitaev model by atoms in an optical cavity have been predicted Mivehvar et al. (2017); Uhrich et al. (2023) and random spin models with atoms in a high-finesse cavity Sauerwein et al. (2023). Quantum simulation of condensed matter phases with a long range photon mediated interaction is a rich topic Caballero-Benitez et al. (2016); Schäfer et al. (2020) including a photon-mediated Peierls transition, emergent quasicrystalline symmetries Mivehvar et al. (2019), multimode BECs Karpov and Piazza (2022), supersolid phases Zhang et al. (2018), spin entanglement and magnetic competition in spinor quantum optical lattices Lozano-Méndez et al. (2022), bond order Caballero-Benitez and Mekhov (2016) as well as other interesting predictions Grankin et al. (2014); Caballero-Benitez and Mekhov (2015); Caballero-Benitez et al. (2016); Caballero-Benitez and Mekhov (2016); Chanda et al. (2022); Mivehvar et al. (2017); Schlawin and Jaksch (2019); Camacho-Guardian et al. (2017); Rylands et al. (2020); Karpov and Piazza (2022); Zhang et al. (2018); Lozano-Méndez et al. (2022). Reviews of cavity QED with quantum gases with varying detail can be found in Refs. Maschler et al. (2008); Mekhov and Ritsch (2012); Ritsch et al. (2013); Mivehvar et al. (2021).
VIII Two-dimensional semi-conductors
Novel semi-conductor materials have emerged as a platform in which new and interesting quantum many-body phases can been realized Kennes et al. (2021). This includes the realization of strongly correlated states such as Wigner crystals Knörzer et al. (2020); Smoleński et al. (2021), Mott insulators Camille Lagoin and Dubin (2022); Erik G. C. P. van Loon and Wehling (2022); Ning Tian (2024), conventional and unconventional superconductivity Yuan Cao and Jarillo-Herrero (2018); Crépel et al. (2022); Han et al. (2023), as well as new forms of hybrid light-matter states Takemura et al. (2014); Sidler et al. (2016); Zhang et al. (2021a); Zhao et al. (2023); Yesenia A. García Jomaso and Pirruccio (2024). Bose-Fermi mixtures can be created by means of electrons interacting with excitons or exciton-polaritons, which to a good approximation are bosonic due to their large binding energy in these 2D materials. In this section, we discuss the recent experimental and theoretical progress regarding mediated interactions between quasiparticles in these materials Tan et al. (2020, 2022); Emmanuele et al. (2020); Song et al. (2022); Muir et al. (2022).
VIII.1 Fermi polaron-polariton interactions
The experimental realization of highly imbalanced Bose-Fermi mixtures in 2D semiconductors allowed for the first observation of Fermi polaron-polaritons Sidler et al. (2016). In these experiments, a two-dimensional electron gas (2DEG) plays the role of the fermionic bath, and excitons play the role of the impurity particles. The excitons can furthermore be hybridised with cavity photons in which case the resulting quasiparticles are referred to as polaron-polaritons.
In a recent experiment using a semiconductor monolayer of MoSe2, the presence of a 2DEG was found to have strong effects on the interaction between polaron-polaritons Tan et al. (2020). A layer of graphene allowed for the tuning of the electron density and the setup was embedded in a high-finesse optical cavity consisting of a distributed-Bragg-reflector to increase the photon-exciton coupling. First, a narrow pump beam tuned to the lower-polariton energy created polaritons, which, due to their excitonic component, interacted with electrons in the 2DEG. The presence of a trion state, i.e., a bound state between an exciton and an itinerant electron, led to an indirect coupling between the polaritons and the trion, forming strongly interacting polaron-polaritons Sidler et al. (2016). A broad probe beam was then applied after a delay time to create and measure the energy of polaron-polaritons in the presence of the polaron-polaritons already created by the pump beam. As seen in Fig. 11 (top), a transient blue shift of the measured polaron-polariton energy was observed for certain delay times between the pump and probe pulses. This blue shift was interpreted as a repulsive interaction between the polariton-polaritons, which was observed to increase significantly when a 2DEG was present as shown in Fig. 11 (bottom). Here the polaron-polariton interactions (purple dots) in the presence of the 2DEG are $\sim 50$ times larger than the bare exciton-polariton interactions (orange dots). Intriguingly, the observed repulsive nature of the polariton-polariton interaction is at odds with the results in Sec. II, which predict the interaction mediated by the electrons to be attractive since the excitons to a good approximation can be regarded as point bosons.
To explain this result theoretically, a Chevy type ansatz Chevy (2006) for the many-body wave function, which has been shown to explain the experimental results in the linear regime Sidler et al. (2016), was employed to study polaron-polaron interactions
$|\Psi_{\mathbf{p}}\rangle=\big{(}\psi^{c}_{\mathbf{p}}\hat{c}_{\mathbf{p}}^{\dagger}+\psi^{c}_{\mathbf{p}}\hat{x}^{\dagger}_{\mathbf{p}}+\sum_{\mathbf{q},\mathbf{k}}\psi_{\mathbf{p},\mathbf{k},\mathbf{p}}\hat{x}^{\dagger}_{\mathbf{p}+\mathbf{q}-\mathbf{k}}\hat{e}^{\dagger}_{\mathbf{k}}\hat{e}_{\mathbf{q}}\big{)}|\text{FS}\rangle.$ (39)
[…]
Writing this many-body wave function formally as $|\Psi_{\mathbf{p}}\rangle=\hat{a}^{\dagger}_{\mathbf{p}}|\text{FS}\rangle$ where $\hat{a}_{\mathbf{p}}^{\dagger}$ creates a polaron-polariton on top of the Fermi sea, the polaron-polariton interactions $U$ was estimated by
$U=\frac{\langle\text{FS}|\hat{a}_{\mathbf{0}}\hat{a}_{\mathbf{0}}\hat{H}\hat{a}^{\dagger}_{\mathbf{0}}\hat{a}^{\dagger}_{\mathbf{0}}|\text{FS}\rangle}{\langle\text{FS}|\hat{a}_{\mathbf{0}}\hat{a}_{\mathbf{0}}\hat{a}^{\dagger}_{\mathbf{0}}\hat{a}^{\dagger}_{\mathbf{0}}|\text{FS}\rangle}-2E_{0}.$ (40)
Here, $E_{0}$ is the energy of a single polaron-polariton and $\hat{H}$ the many-body Hamiltonian. Following this approach, a repulsive interaction was theoretically obtained and attributed to phase-space filling effects.
Alternatively, a theoretical analysis using equilibrium Green’s functions closely following the approach leading to the mediated interaction in Fig. 2(b), predicted an attractive interaction between the polaron-polaritons mediated by the electrons, which is consistent with the results of Sec. III Bastarrachea-Magnani et al. (2021a, b). A blue shift of the polaron-polariton energy could instead be explained by repulsive interactions with trions.
Figure 11: Figure from Ref. Tan et al. (2020). (Top) Probe transmission spectrum for a typical pump-probe experiment. A narrow pump field injects resonantly polaron-polaritons whereas a delayed and broad probe field covers both polariton branches. Yellow dashed lines illustrate the upper and lower polaron-polariton resonances as a function of time delay. (Bottom) Energy shift of the the exciton- and lower polaron-polariton as a function of the concentration of the lower polaron-polaritons.
Strong non-linearities of polaritons immersed in an electron gas were observed in an experiment based on a MoSe2 monolayer in an optical cavity Emmanuele et al. (2020). The interactions between the electrons and the excitons were interpreted in terms of the formation of trion-polaritons instead of the formation of polaron-polaritons, although the coupling of trions to cavity photons is small. However, in the regime where the binding energy of the trion is much larger than the Fermi energy, these two interpretations yield almost the same predictions Glazov (2020). Upon increasing the density of the excitons, two main effects appeared: A quench of the polaron-polariton oscillator strength and positive and negative energy shifts of the lower and middle polaron-polaritons $10-100$ times larger than for neutral exciton-polaritons. These observations were theoretically interpreted as a consequence of the underlying composite electron-hole nature of the excitons, which led to phase-space filling effects. The exchange interaction between two trion-polaritons was predicted to be attractive and $\sim 5$ times stronger than the interaction between exciton-polaritons Song et al. (2022). In Ref. Bastarrachea-Magnani et al. (2021a), it was shown that equilibrium field theory provides an alternative interpretation of the signs as well as magnitudes of the observed energy shifts in terms of interactions with trions.
Using multi-dimensional coherent spectroscopy on a monolayer WS2 Muir et al. (2022), interactions between excitons immersed in a 2DEG were studied at time scales shorter than the typical formation time of Fermi polarons. The observed interaction effects were attributed to phase-filling effects where excitons compete for the same electrons to form polarons. In addition, in Ref. Muir et al. (2022) a bipolaron involving excitons in different valleys with a rather large binding energy was found.
VIII.2 Bose polaron-polariton interactions
Bose polaron-polaritons have also been created in 2D semiconductors using mixtures of exciton-polaritons in two spin states Takemura et al. (2014, 2017); Navadeh-Toupchi et al. (2019). The presence of a bi-exciton state formed by two excitons with opposite spins allowed for the demonstration of a polaritonic Feshbach resonance, leading to the formation of strongly interacting Bose polaron-polaritons. Theoretical predictions showed good agreement with these initial experiments Takemura et al. (2016); Levinsen et al. (2019); Bastarrachea-Magnani et al. (2019). Recently, the interaction between such Bose polarons was explored experimentally in a monolayer MoSe2 embedded in an optical cavity Tan et al. (2022). Using the spin-valley optical selection rules to create a bosonic bath of spin-polarized polaritons mixed with polaritons with opposite spin and a much smaller concentration, the energy of the resulting Bose polarons was measured as a function of their concentration. From this, a polaron-polaron interaction was extracted exhibiting a linear dependence on the density of the bath as shown in Fig. 12(left), which indicates that it is a mediated interaction. The mediated interaction was observed to be attractive for the attractive polaron-polariton and repulsive for the repulsive polaron-polariton. A theoretical approach based on a variational ansatz similar to Eq. 39 was shown to capture the sign-flip of the polariton interactions but predicted an interaction ten times smaller than the experimentally observed, see Fig. 12 (right) Tan et al. (2022).
Figure 12: Figure from Ref. Tan et al. (2022). Left panel: Experimental interaction $g_{ii}$ between attractive (red dots) and repulsive (blue dots) polaron-polaritons as a function of the density $n_{b}$ of the majority excitons. Right panel: Theoretical results from a theory similar to Eq. (40).
These experiments using 2D quantum materials have interesting results whose relation to the interaction between quasiparticles in equilibrium remains to be clarified. The understanding of these systems is typically more complicated than their atomic gas counterparts because the optical excitations can be far from equilibrium, the microscopic interactions between excitons and electrons are not known precisely and not on a simple short range form, and the excitons are composite objects and thus not strictly bosonic particles. This makes a quantitative understanding of the results including the links to Landau’s quasiparticle interaction an open and challenging problem.
IX Mediated collective phases
In this perspective, we have discussed the mediated interaction between individual quasiparticles. When the concentration of these quasiparticles is non-zero, such interactions can give rise to new phases and while this is an extensive topic of its own beyond the scope of this article, we here briefly outline some the recent results in the fields of atomic gases and 2D semiconductors.
Boson-mediated $s$-wave superfluidity.- Many studies have focused on Bose-Fermi mixtures and $s$-wave superfluidity via the exchange of phonon modes. The basic idea is that the bosons form a weakly interacting BEC, which mediates an attractive interaction between two distinguishable fermion species (spin $\uparrow$,$\downarrow$ , ↓). Neglecting retardation effects and assuming weak short range interactions, one can derive analytical results using BCS-theory. In particular, the critical temperature is $k_{B}T_{c}=\frac{\gamma}{\pi}\left(\frac{2}{e}\right)^{7/3}\epsilon_{F}e^{1/\lambda}$ with $\lambda=\frac{m_{c}k_{F}}{2\pi^{2}\hbar^{2}}g_{\uparrow\downarrow}(1-g^{2}/g_{bb}g_{\uparrow\downarrow})$, where $g_{\uparrow\downarrow}$ is the direct interaction between the $\uparrow$ and $\downarrow$ fermions, $g$ is the boson-fermion interaction, and $g_{bb}$ is the boson-boson interaction. Viverit (2002). This result stresses that the superfluid phase can be stabilised by increasing the fermion-boson coupling strength $g$ or by softening the BEC with a smaller $g_{bb}$.
Boson-mediated $p$-wave superfluidity.- In a mixture of single spin fermions and bosons, there cannot be fermionic $s$-wave pairing. The interaction mediated by the bosons can instead induce $p$-wave pairing, which was investigated in early papers Efremov and Viverit (2002); Suzuki et al. (2008). This was extended to strong interactions using Eliashberg theory Wang (2006); Kinnunen et al. (2018); Matsyshyn et al. (2018), as well as to imbalanced Fermi mixtures Wang and He (2019); Wu et al. (2020); Taglieber et al. (2008). Topological $p$-wave pairing in a 2D gas of fermions mixed with a 3D weakly interacting BEC was analysed using strong coupling Eliashberg theory combined with Kosterlitz-Thouless theory Wu and Bruun (2016). In Refs. Bulgac et al. (2006); Bulgac and Yoon (2009); Patton and Sheehy (2012), $p$-wave pairing was predicted to be favorable in spin-imbalanced Fermi mixtures where a fraction of the mixture forms a BEC of strongly bound dimers mediating an attractive interaction between in the remaining spin polarised fermions.
Mediated superfluidity in optical lattices.- There has been extensive interest in exploring induced superfluidity of atoms in optical lattices. In Ref. Illuminati and Albus (2004), the pairing of a two-component Fermi gas mixed with bosons in a 3D optical lattice was studied, predicting an increased critical temperature compared to that coming from an on-site direct interaction. In 2D optical lattices, the long-ranged character of the induced interaction mediated by weakly interacting bosons can compete against a short-ranged direct interaction between fermions, which leads to a rich phase diagram, including $s$-, $p$-, and $d$-wave pairing together with charge and spin density ordering Wang et al. (2005); Klironomos and Tsai (2007). Exotic superfluid phases in 2D optical lattices have been explored in the context of tunable boson-assisted finite-range interactions and Majorana corner modes Wu et al. (2023), topological superfluids Midtgaard et al. (2016); Zhang et al. (2021b), time-reversal-invariant topological superfluids Midtgaard et al. (2017), pairing mediated by hard-core bosons Santiago-García and Camacho-Guardian (2023), FFLO phases Singh and Orso (2020), $p$-wave pairing Wu et al. (2020), and mixed dimensional mixtures Okamoto et al. (2017)
Fermion-mediated collective phases.- The effects of spin-dependent fermion-mediated interactions on the properties of BECs such as stability, phase separation, and depletion were investigated in Refs. Liao (2020); Zheng et al. (2021). In Refs. Huang (2020); Jiang (2021), the intrinsic character of fermion-mediated interactions was explored, and exotic quantum many-body phases resulting from such interactions in optical lattices were proposed Büchler and Blatter (2003); Argüello-Luengo et al. (2022); De and Spielman (2014). The interaction mediated by a superfluid two-component Fermi gas was shown to give rise to strong effects on the sound mode spectrum of a BEC, including avoided crossing features when the sound mode frequencies of the two superfluid match Kinnunen and Bruun (2015). In Ref. Shen et al. (2024), the effects of a fermion mediated interaction on the phase diagram of a Bose-Fermi mixture were explored, and it was shown that strong coupling effects stabilise the mixture again collapse and phase separature well beyond what is expected from mean-field theory. These finding were in good agreement with recent experimental findings Patel et al. (2023).
Exciton-polariton mediated superconductivity.- The experimental realization of exciton-polariton BECs (or coherent states) has motivated several investigations concerning how they can mediate an attractive interaction between electrons thereby inducing superconductivity in a solid state setting. Contrary to atomic gases where the interaction between the fermions can be suppressed, the mediated interaction between electrons has to overcome their Coulomb repulsion. For low condensate densities, the net interaction is typically dominated by Coulomb repulsion, and electron pairing cannot take place. At high densities however, the attractive mediated interaction becomes strong enough to support Cooper pairs and superconductivity. The first proposals studied a 2DEG sandwiched between quantum wells in which a condensate of exctiton-polaritons can mediate an attractive interaction Laussy et al. (2010, 2012); Cotleţ et al. (2016). With the arrival of TMDs, Bogoliubov mode mediated superconductivity recieved renewed attention Skopelitis et al. (2018); Sun et al. (2021a); Yang et al. (2021). This includes the development of a strong-coupling theory of condensate-mediated superconductivity Sun et al. (2021b), and the demonstration of topological superconductivity with a high critical temperature due to the light mass of the polaritons forming the condensate reducing retardation effect Julku et al. (2022). Topological $p$-wave as well as $s$-wave superconductivity mediated by excitons interacting with electrons via a trion state was analysed in Refs. Zerba et al. (2023); von Milczewski et al. (2023). Bound states of two photons have also been predicted to emerge as a consequence of an attractive interaction mediated by a polariton condensate Camacho-Guardian et al. (2021). Recently, spinor BECs have been proposed to enhance superconductivity in TMD materials and atomic systems Bighin et al. (2022).
In the opposite regime where the electron gas can be regarded as the majority compared to the excitons, it has been shown that they can induce a roton minimum in the dispersion of the exciton condensate Shelykh et al. (2010). The presence of a roton mode in ultracold quantum gases has led to the observation of dipolar supersolids Chomaz et al. (2018, 2019); Norcia et al. (2021), and similar effects have been suggested to exist for exciton condensates Matuszewski et al. (2012). The link between fermion-mediated supersolidity and boson-mediated superconductivity has been studied highlighting the importance of the renormalized interaction between polaritons and electrons Cotleţ et al. (2016); Plyashechnik et al. (2023). A recent review on the topic of exciton-photon interactions in 2D microcavities highlights the interplay between many-body physics and photonics Shang et al. (2023). Further details regarding atomic Bose-Fermi mixtures can be found in a recent excellent review Baroni et al. (2024b).
X Conclusions and Outlook
In this perspective article, we discussed how mediated interactions are an intrinsic feature of quasiparticles and explained different theorerical approaches for calculating them. We then described how the advent highly flexible ultracold atomic gases, optical cavities, and 2D TMDs has improved our theoretical as well as experimental understanding of such mediated interactions, and how they open ways to create various interesting quantum phases.
We end by discussing the many interesting and open questions raised by these results. For instance, while the quasiparticle interaction between atomic Fermi polarons has been observed and is well understood for weak to moderately strong interactions, the experimental results for strong interactions are only partly explained by theory as discussed in Sec. VI. Moreover, the quasiparticle interaction between atomic Bose polarons has not been observed so far, which is somewhat surprising since a BEC is more compressible and therefore in general gives rise to a stronger mediated interaction as we saw in Sec. VII. One possible reason is the rather short lifetime of the Bose polaron, which may wash out the effects of a quasiparticle interaction. The interplay between mediated interactions, few-body physics, and the formation of bound states is also an interesting topic that likely will unveil a new panorama to realise exotic many-body phases. So far, studies of mediated interactions have furthermore mostly focused simple environments such as an ideal Fermi gas or a weakly interacting BEC, whereas mediated interactions in strongly correlated phases remain relatively unexplored. Understanding the puzzling results regarding interactions between excitons in TMDs mediated by electrons or other excitons is also a largely open and important problem. This is complicated by several factors such as the composite nature of the excitons leading to phase space filling effects, the non-trival nature of the exciton-electron and exciton-exciton interactions, and the inherent non-equilibrium nature of these systems, which may require the development of new theoretical frameworks. Understanding this can lead to the realisation of new quantum phases in a solid-state setting with far reaching perspectives. All this illustrates that mediated interactions are a crucial concept in quantum many-body physics, and there remains a vast landscape of unexplored possibilities, which could pave the way for breakthroughs in science and technology.
Acknowledgments.- R. P acknowledges Grant No. IN117623 from DGAPA (UNAM) A. C. G. acknowledges financial support from Grant UNAM DGAPA PAPIIT No. IA101923, PAPIME No. PE101223 and PIIF 2023. G. B acknowledges the Danish National Research Foundation through the Center of Excellence CCQ (Grant no. DNRF156).